# Read Anywhere Pointed: Layout-aware GUI Screen Reading with Tree-of-Lens Grounding

### v1

@@ 7 Limitations
In this work, we propose a novel Tree-of-Lens (ToL) agent for the Screen Point-and-Read (ScreenPR) task. Leveraging GPT4-o to generate content and layout descriptions allows us to set up the pipeline quickly and validate ToL’s effectiveness in a short time period. However, the accumulated server delay in requesting GPT4-o services has occupied more than 80% of end-to-end evaluation, not to mention the high expense accompanying, which is a non-trivial issue for real-world applications. Thus, a more efficient and locally hosted model has to be given out in place of GPT-4o. Moreover, since our ToL agent leverages MLLM to generate descriptions, there are potential risks that the generated descriptions from the MLLM could contain harmful contents. Therefore more post-processing strategy might be needed. Last but not least, another critical challenge is balancing application efficiency with improved semantic representation, thereby enabling on-device support and further expanding ToL’s application areas. Our current work is limited for research purpose and more application-wise exploration is pending.
@@ Appendix A Example outputs of the ToL Agent
For each GUI domain, we show one example of the output of our ToL agent along with the lenses generated. We do not deliberately cherry pick the best case, but show examples with both strong and weak performances. An example for the operation system domain is shown in Figure 6. An example for the mobile domain is shown in Figure 7. An example for the web domain is shown in Figure 8.
@@ Appendix B Examples of the Cycle Consistency Evaluation
We select one result of our cycle consistency evaluation for the content description and show it in Figure 9. The left four images come from the cropped areas from different GUI screens, with only one of which matches the description in the prompt on the right side. As shown, GPT4-o can successfully choose the correct option in the "Answer" field and give its reasoning in the "Analysis" field.
@@
{coloredquote}
Given the following information in a mobile navigation task:
Historical action and region description: {{haction}}
Task Goal: {{goal}}
Current region: {{region}}
Please provide your answer in the following JSON format:
"Analysis": "…",
"Answer": "yes/no"
Figure 12: Prompt used by GPT-4o in Action Validation
@@
Appendix C Text Prompt templates
@@
Appendix D Mobile Navigation Action Verification
D.1 Example
@@
D.2 Sample Strategy of Mobile Navigation Action Execution Trajectory
@@
D.3 Action verification Pipeline Details
As shown in Figure 5, a set of navigation actions, target points, and corresponding screenshots, are first recorded during the interaction between the mobile navigation agent and the mobile screen. Then, our ToL agent first processes each pair of target points and screenshots $(P_{i},S_{i})$, including those from the history path of navigation trajectory and the currently planned one $(i\in[0,t-1])$, where the $P_{t}-1$ is the target point that the navigation agent plans to interact with for the next step. As a result, the layout and content descriptions $\hat{D}^{c}_{i}$, $\hat{D}^{l}_{i}$ are generated. Next, we concatenate all the descriptions and send them to the GPT-4-Turbo model to determine whether agent actions follow the given task goal.
@@ D.3 Action verification Pipeline Details
Appendix E Human Evaluation
We leverage human evaluation to test different screen reading models on our ScreenPR benchmark. We leverage the Amazon Mechanical Turk [^4] https://www.mturk.com platform to find annotators to complete the survey that we designed for our human evaluation. We pay \$0.15 for each survey and the screenshot of the survey is shown in Figure 18
Appendix F GUI Region Detection Model

### v2

@@ 7 Limitations
In this work, we propose a novel Tree-of-Lens (ToL) agent for the Screen Point-and-Read (ScreenPR) task. By leveraging GPT4-o to generate content and layout descriptions, we are able to set up the pipeline quickly and validate the effectiveness of ToL grounding method in a short time period. However, the accumulated server delay in requesting GPT4-o services has occupied more than 80% of end-to-end evaluation, not to mention the high expense accompanying, which is a non-trivial issue for real-world applications. Thus, a more efficient and locally hosted model has to be given out in place of GPT-4o. Moreover, since our ToL agent relies on a certain MLLM to generate descriptions, there are potential risks that the generated descriptions from the MLLM could contain harmful contents. Therefore more post-processing strategy might be needed. Last but not least, another critical challenge is balancing application efficiency with improved semantic representation, thereby enabling on-device support and further expanding ToL’s application areas. Our current work is limited for research purpose and more application-wise exploration is pending.
@@ Appendix A Example outputs of the ToL Agent
For each GUI domain, we show one example of the output of our ToL agent along with the lenses generated. We do not deliberately cherry-pick the best case but show examples with both strong and weak performances. An example for the operation system domain is shown in Figure 6. An example for the mobile domain is shown in Figure 7. An example for the web domain is shown in Figure 8.
@@ Appendix B Examples of the Cycle Consistency Evaluation
We select one result of our cycle consistency evaluation for the content description and show it in Figure 9. The left four images come from the cropped areas from different GUI screens, with only one of which matches the description in the prompt on the right side. As shown, GPT-4o can successfully choose the correct option in the "Answer" field and give its reasoning in the "Analysis" field.
@@
{coloredquote}
Given the following information in a mobile navigation task:
Historical action and region description: {{haction}}
Task Goal: {{goal}}
Current region: {{region}}
Please provide your answer in the following JSON format:
"Analysis": "…",
"Answer": "yes/no"
Figure 15: Prompt used by GPT-4o in Action Validation
@@
Appendix C Cross-evaluation Experiments with Cycle Consistency Evaluation
Our cycle consistency evaluation, relying on GPT-4V as the auxiliary model, shows that our ToL agent (leveraging GPT-4o by default) outperforms the other baselines including the vanilla GPT-4o. Since the evaluator, the baseline model, and our ToL agent all rely on GPT models, we believe the results regarding models relying on GPT models are less influenced by the potential bias as mentioned by the review. Furthermore, to strengthen our result, we first include the Gemini Pro Vision model as a baseline model and introduce a new version of our ToL agent relying on the Gemini Pro Vision model. Then, we extend our main results with the cross-evaluation setting where the models relying on GPT-4o will not be evaluated with GPT-4V as the auxiliary model, and the models relying on Gemini Pro Vision will not be evaluated with Gemini Pro Vision as the auxiliary model. The results are shown in Table 8, where under this cross-evaluation setting, our ToL agent consistently demonstrates clear superiority over the vanilla MLLM baseline.
Auxiliary Model: $\begin{array}[]{l}\text{Auxiliary Model: }\\ \text{GPT-4V}\end{array}$
Content Acc Layout Acc
Gemini Pro Vision 33.20 12.27
ToL Agent (Gemini Pro Vision) 45.33 18.53
Auxiliary Model: Gemini Pro $\begin{array}[]{l}\text{Auxiliary Model: }\\ \text{Gemini Pro Vision}\end{array}$
Content Acc Layout Acc
GPT-4o 43.52 22.33
ToL Agent (GPT-4o) 58.44 33.93
Table 8: Cycle consistency evaluation of the ToL agent and vanilla MLLMs under the cross-evaluation settings. The performance of the ToL agent and the baseline MLLM are evaluated with the auxiliary model being a different MLLM than the one under test. The ToL agent consistently outperforms the baselines.
Appendix D Text Prompt templates
@@
Appendix E Mobile Navigation Action Verification
E.1 Example
@@
E.2 Sample Strategy of Mobile Navigation Action Execution Trajectory
@@
E.3 Action verification Pipeline Details
As shown in Figure 5, a set of navigation actions, target points, and corresponding screenshots, are first recorded during the interaction between the mobile navigation agent and the mobile screen. Then, our ToL agent first processes each pair of target points and screenshots $(P_{i},S_{i})$, including those from the history path of navigation trajectory and the currently planned one $(i\in[0,t-1])$, where the $P_{t}-1$ is the target point that the navigation agent plans to interact with for the next step. As a result, the layout and content descriptions $\hat{D}^{c}_{i}$, $\hat{D}^{l}_{i}$ are generated. Next, we concatenate all the descriptions and send them to the GPT-4 to determine whether agent actions follow the given task goal.
@@ E.3 Action verification Pipeline Details
Appendix F Human Evaluation
We leverage human evaluation to test the performance of different screen readers on our ScreenPR benchmark. We leverage the Amazon Mechanical Turk [^4] https://www.mturk.com platform to find annotators to complete the survey that we designed for our human evaluation. We hired 3 human judges for each data sample. The judges are asked to rate the generated descriptions based on three criteria: how well they describe the content, how well they articulate the layout, and their overall preference. Each question has four options, ranging from "very well", “fair”, “not well”, and "awful," which we numerically mapped to accuracy scores of 100%, 66%, 33%, and 0%, respectively. We pay \$0.15 for each survey and the screenshot of the survey is shown in Figure 18. Approximately 81% of the data samples receive at least two identical ratings from the human judges, and the standard deviation is 19.4%.
Appendix G GUI Region Detection Model